# Trapped fireshell (halo) of photons and pairs around black-hole horizon: source for ultra-high-energy particles

### v1

@@ 2.2 Thermodynamic and opaque fluid of photons and pairs
The local temperature $T_{\gamma}(r)$ represents the mean kinematic energy of photons and electron-positron pairs, which possess zero chemical potential. As a result, the trapped fireshell is an electrically neutral, opaque and thermalised fluid of photons and electron-positron pairs. We describe it by a thermodynamic state of thermal temperature $T_{\gamma}(r)$, energy and number densities
@@
in the local frame at the radius $r$. The Stefan-Boltzmann constant $\sigma_{{}_{\rm SB}}=\pi^{2}/15$ for setting the Boltzmann constant $k_{B}=1$. Here, we adopt $d_{f}\approx 30$ for the number of massless gauge bosons and light fermion-antifermion pairs in the Standard Model (SM) of fundamental particle physics  [^1] Since the number density is so high, see Fig. 3, neutrinos and anti-neutrinos can be possibly kept inside the fluid in participation of thermalisation., which can kinematically participate in the thermal equilibration at temperature $T_{\gamma}>10$ MeV. The boundary conditions (2.4) gives the horizon temperature $T_{\gamma}^{+}$ near the horizon $r_{+}=2GM$,
@@
whose value depends on the balance between ultra-relativistic particles and gravitational pressures. In general, the horizon temperature represents the maximum temperature of trapped photon-pair spots at the place where the gravitational force is maximum. The horizon temperature $T^{+}_{\gamma}>2m_{e}$ is a necessary condition, and $T_{\gamma}^{+}>10$ MeV for the photon and pair energy density $\rho_{\gamma}^{+}\approx\rho_{n}$ is about the nuclear density. We use the electron-positron pair ($e^{-}e^{+}$) notion for all possible SM pairs, and use the subscript $\gamma$” for the quantities of the trapped fireshell of photons and pairs.
@@
and $n_{\gamma}^{+}\equiv n_{\gamma}(r_{+})$, corresponding to the total energy $E_{\gamma}$ (2.5). Here, the macroscopic length $\delta\ell$ is approximately equal to the one in Eq. (2.5) for the total energy. Therefore, we effectively describe the trapped fireshells by the temperature $T^{+}_{\gamma}$ and length $\delta\ell$.
@@
As shown in Fig. 3, the mean-free path (2.7) are small; photons and pairs are not free streaming outward, and the luminosity $L_{\gamma}=4\pi r^{2}\rho_{\gamma}c$ represents the diffusion flux of photons and pairs inside the halo. In the microscopic scale around each local point $r$ of the halo, random particle collisions in all directions, the thermal equilibrium state and temperature $T_{\gamma}(r)$ are isotropic without any preferential direction. However, at the macroscopic scale, the temperature $T_{\gamma}(r)$ gradient is large in the radial direction due to the strong gravitational field near the horizon.
@@
We use the local temperature $T_{\gamma}$ to quantitatively describe a macroscopic-sized halo of the opaque thermal fluid of energetic photons and pairs, which is trapped around spherical black hole horizons following massive core collapses. One can generalise such a halo to macroscopic-sized spots in strong gravity and non-spherically symmetric environments of astrophysical accretion and coalescence in the presence of angular momentum, magnetic fields and other components.
Such a temperature $T_{\gamma}(r)$ gradient results in a radially outward diffusion flux of photons and pairs, dissipating the halo energy $E_{\gamma}$ (2.5) to the blackbody radiation and the emissions of UHE particles, as speculated in Refs. [24, 25] and will be discussed below. Therefore, the halo should be a metastable configuration that gradually shrinks and disappears. However, compared with the halo hydrodynamical relaxation On a time scale $2r_{+}/v_{s}$, the rate of energy dissipation is low, so we approximate the halo as a steady-state configuration of thermal energy sources.
We now turn to discussions of the dynamics of accelerating UHE particles on microscopic scales $\sim 10^{-2}$ cm. The Such a dynamics depends only on the local temperature of the spots, independently of the macroscopic sizes and global symmetries of the fire spots.
@@ 3 Compton-rocket effect on accelerating electrons
3.1 Optically thin plasma of electrons and protons irradiated by photons
Reference [5] studies the Compton-rocket effect: via Thomson scattering, the radially outward photon flux (intensity) $F_{\gamma}(r)=\rho_{\gamma}(r)c$ acts as a radiation force on an optically thin gas of hot electrons or cold protons, and accelerates them to high energies. The photon-scattering excess force on an isotropic gas results from the globally anisotropic loss of internal radiation energy. The rates of energy and momentum transfer from the radiation flux $F_{\gamma}(r)$ to the electron gas are (per electron),
@@ 3.1 Optically thin plasma of electrons and protons irradiated by photons
with the accelerated electron velocity $\beta_{e}=v_{e}/c$, Lorentz factor $\gamma_{e}=(1-\beta^{2}_{e})^{1/2}$, $(\beta_{e}\gamma_{e})^{2}=\gamma_{e}^{2}-1$ and momentum $p_{e}=m_{e}\beta_{e}\gamma_{e}$. Here, we use a local observer’s proper time $\tau$ in a flat Lorentz frame at the radius $r$. Moreover, to simplify the notations consistently with the following discussed one-dimensional model, we do not specify the averages over the electron distribution in the momentum phase space [5]. Equation (3.2) gives the Eddington luminosity argument for the incident radiation pressure on non-relativistic electrons ($\beta_{e}\approx 0$) in the Thomson scattering limit. For the proton case, these formulae remain the same with the substitutions of the subscript $e\rightarrow p$ and approximate cross section $\sigma_{{}_{T}}\rightarrow\sigma^{p}_{{}_{T}}\approx(m_{e}/m_{p})^{2}\sigma_{{}_{T}}\ll\sigma_{{}_{T}}$ of multi-photons and proton collisions.
@@
It is worthwhile to emphasise that the Compton-rocket effect (3.1,3.2), where present the photon flux $F_{\gamma}(r)$, is essentially attributed to the nonlinear QED processes of multi-photon scatterings and interactions (see, e.g., [26, 27, 28]). These non-linear interactions are particularly important when the radiation field is energetic and the charged particle density is large. It is what we will study in the following sections.
3.2 Electrons and protons in an opaque fluid of photons and pairs
The trapped fireshell (halo) of photons and pairs around the black hole horizon has an ultra-high temperature $T_{\gamma}$, see Fig. 3. All material electrons and protons are ionised inside the fluid. Their average densities are much smaller than the density of photons and pairs. They have nonzero chemical potentials and do not participate in the thermal equilibrium of photons and pairs via the process $\gamma+\gamma\leftrightarrow e^{+}e^{-}$. They couple with photons and pairs via multi-particle collisions and field interactions, and are in an energy equipartition with photons and pairs. Electrons are relativistic, protons are non-relativistic, and they do not form an electron-proton oscillating plasma. We describe them by the number of bulk electrons and protons inside the fluid,
$\bar{N}_{e}$ $=$ $\bar{N}_{p}=BN_{\gamma},~~~\gamma_{e}=\langle\tilde{\gamma}_{e}\rangle=1,~~~\beta_{e}=\langle\tilde{\beta}_{e}\rangle=0.$ (3.3)
Here, the baryon-loading $B\ll 1$ quantifies the total number of electrons (protons) inside an opaque spot of photons and pairs of the number $N_{\gamma}$. The $B$ should relate to the baryon loading of fireballs, whose outward expansion accounts for GRBs. The $\tilde{\gamma}_{e}$ and $\tilde{\beta}_{e}$ represent the velocity and Lorentz factor of bulk electrons, randomly colliding with photons and pairs. The averages $\langle\tilde{\beta}_{e}\rangle=0$ and $\langle\tilde{\gamma}_{e}\rangle=1$ are over the distribution of bulk electrons.
@@ 3.2 Electrons and protons in an opaque fluid of photons and pairs
In a one-dimensional model for the spherical symmetric fireshell, the bulk electrons averaged momentum $\langle\tilde{p}_{e}\rangle=m_{e}\langle\tilde{\beta}_{e}\gamma_{e}\rangle$ vanishes with the random distribution $\propto e^{-\tilde{p}_{e}^{2}/T_{\gamma}}$ with an equal probability of inward and outward moving. Thus, bulk electrons do not have a net drift in the direction $\hat{r}>0$ or opposite one $\hat{r}<0$, and their averaged velocity vanishes $\langle\tilde{\beta}_{e}\rangle=0$ and $\langle\tilde{\gamma}_{e}\rangle=1$, and the same discussions for bulk protons $\bar{N}_{p}$.
@@
However, the situation changes when the strong radiation force (3.2) is present, acting on bulk electrons, since the density of photons and pairs has a large variation gradient in the radial direction, as shown in Figs. 2 and 3. We attempt to investigate the Compton-rocket effect on bulk electrons in a highly opaque fluid of photons and pairs at high temperatures $T_{\gamma}\gg m_{e}$, densities $\rho_{\gamma}\gg m_{e}^{4}$ and $n_{\gamma}\gg m_{e}^{3}$. Namely, how the radially outward flux (intensity) $F_{\gamma}(r)=\rho_{\gamma}(r)c\gg m_{e}^{4}$ acts as a radiation force on bulk electrons and protons (3.3), accelerating them outward along the radial direction to high energies.
@@
3.3 Drifting electrons in an opaque fluid of photons and pairs
Although the mean-field path of bulk electrons is small, as shown in Fig. 3, the outward radiation force $F_{\gamma}=\rho_{\gamma}(r)c$ is strong, so some bulk electrons have chances within their mean-free path to get accelerated and drift radially outwards. Therefore, some bulk electrons drift in the direction $\hat{r}>0$ and their averaged momentum $p_{e}=\langle\tilde{p}_{e}\rangle_{a}$ and energy $\epsilon_{e}=(p_{e}^{2}+m_{e}^{2})^{1/2}$ do not vanish. We define the number density of these outward drifting electrons as

### v2

@@ 2.3 Thermodynamic and opaque radiation fluid of photons and pairs
The local temperature $T_{\gamma}(r)$ represents the mean kinematic energy of photons and electron-positron pairs, which possess zero chemical potential. As a result, the trapped fireshell is an electrically neutral, opaque and thermalised fluid of photons and electron-positron pairs, which we also call a radiation fluid for short. We describe it by a thermodynamic state of thermal temperature $T_{\gamma}(r)$, energy and number densities
@@
in the local frame at the radius $r$. The Stefan-Boltzmann constant $\sigma_{{}_{\rm SB}}=\pi^{2}/15$ for setting the Boltzmann constant $k_{B}=1$. The temperature $T_{\gamma}(r)$ (2.9), Fig. 2 (right), is the averaged kinetic energy of relativistic particles over species $d_{f}$. In the Standard Model (SM) of fundamental particle physics, we adopt $d_{f}\approx 30$ for the number of massless gauge bosons and light fermion-antifermion pairs  [^5] Since the number density is so high, see Fig. 3, neutrinos and anti-neutrinos can be possibly kept inside the fluid in participation of thermalisation., which can kinematically participate in the thermal equilibration at temperature $T_{\gamma}>10$ MeV. The boundary conditions (2.5) gives the temperature $T_{\gamma}^{+}$ near the horizon,
@@
whose value depends on the balance between ultra-relativistic particles and gravitational pressures. The horizon temperature $T^{+}_{\gamma}$ characterises the maximum temperature of a trapped photon-pair fireshell at the place where the gravitational force is maximum. For these discussions, the equipartition of particle species’ kinetic energies is necessary; an exact thermal equilibrium is not required.
@@
corresponding to the total energy $E_{\gamma}$ (2.6). Here, the maximal photon-pair number density near the horizon is $n_{\gamma}^{+}\equiv n_{\gamma}(r_{+})$, the macroscopic length $\delta\ell$ is approximately equal to the one in Eq. (2.6) for the total energy. Therefore, we effectively describe the trapped fireshells by the temperature $T^{+}_{\gamma}$ and length $\delta\ell$.
The proper time and energy of physical processes occurring in the fireshell near the horizon receive significant gravitational dilation and redshift with respect to an observer at infinity, see Fig. 1 and its caption. The fireshell differs from the photon sphere (ring) of radius $3r_{+}/2$, which describes photons (light) forced to travel in orbits, due to the strong gravitational field near the horizon.
2.4 Temperature radial gradient and anisotropic radiation flux
@@
Photons and pairs in the fireshell are not freely streaming outward, because the mean-free path (2.8) is small, shown in Fig. 3. On the microscopic scale, a local spacetime around each local point $(r,t)$ inside the fireshell, random particle collisions in all directions, the thermal equilibrium state and temperature $T_{\gamma}(r)$ are spatially isotropic without any preferential direction. However, on the macroscopic scale, the temperature $T_{\gamma}(r)$ has a large gradient in the radial direction due to the strong gravitational field near the horizon. Such a temperature $T_{\gamma}(r)$ gradient results in a radially outward diffusion of photons and pairs inside the fireshell. Following the photon-pair energy density $\rho_{\gamma}(r)$ and its value (2.5) near the horizon, we describe the diffusion by the flux $F_{\gamma}=\rho_{\gamma}c$ and luminosity $L_{\gamma}=4\pi r^{2}\rho_{\gamma}c$, whose values near the horizon are
$F_{\gamma}^{+}\equiv\rho_{\gamma}(r_{+})c=\rho_{n}c,~~~L_{\gamma}^{+}\equiv L_{\gamma}(r_{+})=\eta L_{\rm pl},$ (2.12)
where the “Planck luminosity” $L_{\rm pl}\equiv 4\pi r_{+}^{2}\rho_{{}_{M}}c=(3/2)m^{2}_{\rm pl}=3.73\times 10^{59}{\rm ergs/sec}$.
As speculated in Refs. [17, 18] and will be discussed below, the diffusion flux dissipates the fireshell energy $E_{\gamma}$ (2.6) to the blackbody radiation and the emissions of UHE particles. Therefore, the fireshell energy density and pressure $P_{\gamma}=\rho_{\gamma}/3$ decrease. The fireshell is a metastable configuration that gradually shrinks due to gravity. Since the fireshell opacity is large, the energy dissipation rate $\dot{E}_{\gamma}/E_{\gamma}$ is small, and processes are nearly adiabatic. Thus, we will approximately consider the fireshell as a steady-state configuration (source) with a large thermal energy $E_{\gamma}$ (2.6) budget and an enormous number of relativistic particles $N_{\gamma}$ (2.11).
In this fireshell example for highly energetic arenas, we use the TOV equation (2.2) and local temperature $T_{\gamma}(r)$ to describe a macroscopic-sized fireshell of the opaque thermal fluid of dense and energetic photons and pairs, which is trapped around a spherical black hole horizon following a massive stellar core collapse. We ought to mention the following points.
  1. (1)
@@
    One can generalise such a thermal fireshell to macroscopic-sized firespots in strong gravity and non-spherically symmetric environments of astrophysical accretion and coalescence in the presence of angular momentum, magnetic fields and other components.
  2. (2)
    The gravity strength decrease results in the energy density $\rho_{\gamma}$ and temperature $T_{\gamma}$ decreases, the latter is also affected by the number $d_{f}$ (2.9) of particle species participating in the thermal equilibrium. When the temperature (mean kinetic energy) drops below the hadronic binding energy, hadronisation occurs, and the photon-lepton(pair) and gluon-quark(pair) thermal configuration becomes the photon-lepton(pair) one with electrons, protons and neutrons.
As will be discussed below, the dynamics producing UHE particles are on microscopic scales $\sim 10^{-2}$ cm, independently of firespots’ macroscopic shapes and global symmetries. The essentially relevant quantities for the dynamics are maximum temperature $T_{\gamma}^{+}>10$ MeV, the macroscopic size $\delta\ell$ (2.6) and baryon matter contamination $B$ of the firespots.
2.5 Baryon matter contamination, bulk electrons and protons
The trapped fireshell of photons and pairs around the black hole horizon has an ultra-high temperature $T_{\gamma}$, see Fig. 3, which, however, drops quickly below $100$ MeV as the radius increases from the horizon. The quark-gluon plasma undergoes a phase transition of hadronisation and becomes a baryon matter, some of which contaminates the radiation fluid of photons and pairs inside the fireshell. We are interested in the temperature range $T_{\gamma}\sim(10\sim 50)$ MeV, and the baryon matter inside the fireshell is fully ionised, comprising electrons ($e$) and protons ($p$), which are electrically neutral in total.
These electrons ($e$) and protons ($p$) have to be distinguished from electrons $e^{-}$ and positrons $e^{+}$ associated with two photons $2\gamma$ (2.1). They have non-zero chemical potentials and do not participate in the back-and-forth process $\gamma+\gamma\leftrightarrow e^{+}+e^{-}$ (2.1) for the thermal equilibrium of photons and pairs. Therefore, their mean kinetic energies do not share with the mean kinetic energies (temperature $T_{\gamma}$) of photons and pairs. These electrons ($e$) and protons ($p$) effects on the thermal equilibrium, which were studied by using the Boltzmann equation in Ref. [15], are small if their average energy and number densities are much smaller than those of photons and pairs.
Thus, we model such an ionised baryon matter contamination as dilute electron ($e$) gas and proton ($p$) gas immersed inside the thermal and neutral radiation fluid of photons and pairs at the temperature $T_{\gamma}$. In this article, we focus on these electrons ($e$) and protons ($p$), their electromagnetic interaction with a dense radiation fluid of photons and pairs. We ignore neutrons and weak interactions associated with the inverse $\beta$ decay and anti-neutrino production.
Colliding with photons and pairs via multi-particle collisions and field interactions, these electrons ($e$) and protons ($p$) are in microscopic random motions in all directions. They form a relativistic ($e$) gas and a non-relativistic ($p$) gas, and do not form an electron-proton oscillating plasma. We describe them by bulk electrons and protons inside the fireshell, whose numbers are,
$\bar{N}_{e}$ $=$ $\bar{N}_{p}=BN_{\gamma}.$ (2.13)
Here, the baryon-loading $B\ll 1$ quantifies the total number of electrons (protons) inside an opaque fireshell of photons and pairs of the number $N_{\gamma}$. The $B$ should relate to the baryon loading of fireballs, whose outward expansion accounts for GRBs.
We adopt the notations $\tilde{\beta}_{e}$ and $\tilde{\gamma}_{e}$ to represent the velocity and Lorentz factor of bulk electrons  [^6] Bulk electrons (protons) are relativistic (non-relativistic), which indicates individual bulk electron (proton) of Lorentz factor $\tilde{\gamma}_{e}>1$ ($\tilde{\gamma}_{p}\gtrsim 1$) without being averaged over their distribution function in energy-momentum phase space. It is analogous to an opaque radiation fluid of relativistic photons and pairs, whose macroscopic diffusion (drifting) velocities are small, corresponding to an averaged Lorentz factor close to unity., which are in random motions at microscopic scales, therefore,
$\langle\tilde{\beta}_{e}\rangle=0,~~~\langle\tilde{\gamma}_{e}\rangle=1,$ (2.14)
where the averages $\langle\tilde{\beta}_{e}\rangle$ and $\langle\tilde{\gamma}_{e}\rangle$ are over the energy-momentum distribution of bulk electrons. The average values (2.14) mean that bulk electrons have no macroscopic motions in any preferential direction. However, due to an external radiation force or other dynamics at macroscopic scales, the average values can be non-zero: $\langle\tilde{\beta}_{e}\rangle\not=0$ and $\langle\tilde{\gamma}_{e}\rangle>1$. It implies that some of the bulk electrons tend to macroscopic motions in a preferential direction and with a non-vanishing velocity
$\beta_{e}=\langle\tilde{\beta}_{e}\rangle\not=0,~~~\gamma_{e}=\langle\tilde{\gamma}_{e}\rangle>1.$ (2.15)
We call them (2.15) drifting electrons with a non-vanishing velocity $\beta_{e}\not=0$ and Lorentz factor $\gamma_{e}>1$ at macroscopic scales. The same discussions apply to bulk and drifting protons. We now turn to discuss the dynamics of radiation force accelerating electrons from the bulk electrons (2.14) to the drifting electrons (2.15).
@@ 3 Compton-rocket effect on accelerating electrons
3.1 Radiation force acting on bulk electrons
Reference [21] studies the Compton-rocket effect: via Thomson scattering, the radially outward photon flux (intensity) $F_{\gamma}(r)=\rho_{\gamma}(r)c$ acts as a radiation force on an optically thin gas of hot electrons or cold protons, and accelerates them to high energies. The photon-scattering excess force on an isotropic gas results from the globally anisotropic loss of internal radiation energy. In the one-dimensional model with respect to a local observer’s proper time $\tau$ in a flat Lorentz frame at the radius $r$, the rates of energy and momentum transfer from the radiation flux $F_{\gamma}(r)$ to the electron gas are (per electron),
@@ 3.1 Radiation force acting on bulk electrons
and some of the bulk electrons [^7] The phrase “some of the bulk electrons” will be quantified by the fraction $N_{e}/\bar{N}_{e}$ (3.4), which equivalently corresponds to the probability of a bulk electron accelerated to the macroscopic motion along the radial direction. are accelerated to drifting electrons with macroscopic velocity $\beta_{e}=v_{e}/c$, Lorentz factor $\gamma_{e}=(1-\beta^{2}_{e})^{1/2}$, and momentum $p_{e}=m_{e}\beta_{e}\gamma_{e}$, $(\beta_{e}\gamma_{e})^{2}=\gamma_{e}^{2}-1$, corresponding to (2.15). The initial conditions at $\tau=0$ for these dynamical equations (3.1) and (3.2) are $\beta_{e}=\langle\tilde{\beta}_{e}\rangle=0$ and $\gamma_{e}=\langle\tilde{\gamma}_{e}\rangle=1$ (2.14), when electrons are bulk electrons. Note that the radiation flux $F_{\gamma}(r)$ due to multi-photon scattering off an electron should be considered as a collective force at macroscopic scales, rather than a single photon scattering off an electron at microscopic scales.
@@
To simplify the notations consistently with the one-dimensional model, we do not specify the averages $\langle\cdot\cdot\cdot\rangle$ over the electron distribution in the energy-momentum phase space. Equation (3.2) gives the Eddington luminosity argument for the incident radiation pressure on non-relativistic electrons ($\beta_{e}\approx 0$) in the Thomson scattering limit [21]. For the proton case, these formulae and discussions remain the same with the substitutions of the subscript $e\rightarrow p$ and approximate cross section $\sigma_{{}_{T}}\rightarrow\sigma^{p}_{{}_{T}}\approx(m_{e}/m_{p})^{2}\sigma_{{}_{T}}\ll\sigma_{{}_{T}}$ of multi-photons and proton collisions.
@@
Radiation fields non-linearly interact with relativistic electrons via the Compton rocket effect [21, 22, 23] by the anisotropy of multi-photons and electron scattering. This dynamics has been studied in an optically thin plasma of electrons and protons by particle-in-cell numerical simulations [24, 25, 26, 27]. It is worthwhile to emphasise that the Compton-rocket effect (3.1,3.2), where the photon flux $F_{\gamma}(r)$ is essentially attributed to the nonlinear QED processes of multi-photon scatterings and interactions (see, e.g., [28, 29, 30]). These non-linear interactions are particularly important when the radiation field is energetic and the charged particle density is large. It is what we will study in the following sections.
3.2 Drifting electrons in an opaque radiation fluid at high temperatures
@@
In the one-dimensional and spherical symmetric fireshell, the bulk electrons averaged momentum $\langle\tilde{p}_{e}\rangle=m_{e}\langle\tilde{\beta}_{e}\tilde{\gamma}_{e}\rangle$ vanishes with the random momentum distribution $\propto e^{-\tilde{p}_{e}^{2}/T_{\gamma}}$ with an equal probability of inward and outward moving. Thus, bulk electrons do not have a net drift motion in the direction $\hat{r}>0$ or opposite one $\hat{r}<0$, and their averaged velocity vanishes $\langle\tilde{\beta}_{e}\rangle=0$ and $\langle\tilde{\gamma}_{e}\rangle=1$ (2.14), and the same discussions for bulk protons $\bar{N}_{p}$.
@@
However, the situation changes when the strong radiation force (3.2) is present, acting on bulk electrons, since the density of photons and pairs has a large variation gradient in the radial direction, as shown in Figs. 2 and 3. We attempt to investigate the Compton-rocket effect on bulk electrons in a highly opaque fluid of photons and pairs at high temperatures $T_{\gamma}\gg m_{e}$, densities $\rho_{\gamma}\gg m_{e}^{4}$ and $n_{\gamma}\gg m_{e}^{3}$. Namely, how the radially outward flux (intensity) $F_{\gamma}(r)=\rho_{\gamma}(r)c\gg m_{e}^{4}$ acts as a radiation force on bulk electrons and protons (2.14), accelerating some of them outward along the radial direction to ultra high energies, drifting out of the fireshell.
@@
3.3 Drifting electrons versus bulk electrons
Although the mean-field path of bulk electrons is small, as shown in Fig. 3, the outward radiation force $F_{\gamma}=\rho_{\gamma}(r)c$ is strong, so some bulk electrons have chances within their mean-free path to get accelerated and drift radially outwards. Therefore, some of bulk electrons drift in the direction $\hat{r}>0$ and their averaged momentum $p_{e}=\langle\tilde{p}_{e}\rangle_{a}$ and energy $\epsilon_{e}=(p_{e}^{2}+m_{e}^{2})^{1/2}$ do not vanish. They are drifting electrons with the macroscopic velocity $\beta_{e}$ and Lorentz factor $\gamma_{e}$ (2.15), and we define the number of these outward drifting electrons as